### v1

@@ Abstract.
Here, for $d\in\{2,3\}$, $\bm{v}$ represents the velocity field, $\pi$ denotes the pressure, and $\bm{f}$ is the external forcing term. The constants $\kappa$ and $\nu$ correspond to the relaxation time and kinematic viscosity, respectively. The parameter $p\in(1,\infty)$ characterizes the fluid’s flow behavior, and $\mathbf{D}(\bm{v})$ denotes the symmetric part of the velocity gradient $\nabla\bm{v}$. For the power-law exponent $p\in\mathopen{}\mathclose{{\left(\frac{2d}{d+2},\infty}}\right)$, we establish the existence of a weak solution to the generalised Navier-Stokes-Voigt equations. Furthermore, we demonstrate that the weak solution is unique for the same range of the exponent $p$. The optimality of our results lies in the framework’s use of a Gelfand triple, which allows the Aubin–Dubinskiĭ lemma to yield strong convergence of approximate solutions, essential for existence and valid precisely for $p>\frac{2d}{d+2}$.
@@ 1.2. Literature review
$1<p<\frac{2d}{d+2},\quad\mbox{for}\ \ d\not=2,$
@@ 1.3. Objectives and novelties of the paper
  • As mentioned above, there are few works in the literature addressing power-law fluids [2, 3, 7, 21, 51]. In [51], the authors studied power-law fluids and established existence results for the exponent $p\in\big(\frac{2(d+1)}{d+2},\infty\big)$ using the $\mathrm{L}^{\infty}$-truncation, while [21] extended these results to $p\in\big(\frac{2d}{d+2},\infty\big)$ via Lipschitz-truncation. In our analysis, the presence of the Voigt term $\kappa\partial_{t}\Delta\bm{v}$ allows us to establish existence and uniqueness of the generalised Navier–Stokes–Voigt equations for the optimal range $p\in\big(\frac{2d}{d+2},\infty\big)$ using compactness and monotonicity arguments.
  • Another novelty of this article is that we do not employ the Helmholtz–Hodge projection, which would restrict our analysis to the case $p\in[2,\infty)$. Although the Helmholtz–Hodge projection is well-defined for $p\in(1,2)$, the analysis in this range is highly delicate and creates significant difficulties in controlling the convective term in the momentum equation (1.1)1. By avoiding the Helmholtz–Hodge projection, we are able to treat the full range $p>\frac{2d}{d+2}$ directly, without encountering these technical obstacles.
@@ 1.4. Organization
The structure of this article is organized as follows: Section 2 lays the groundwork by introducing the necessary function spaces and auxiliary results essential for the analysis. In Section 3, we present the definition of the weak solution (Definition 3.1) and state our main results, including Theorems 3.2 and 3.3. As previously mentioned, the proof of Theorem 3.2 is divided into two cases based on the range of the exponent $p$ (see (3.6)). Section 4 focuses on the proof of Theorem 3.2 for the case of large $p$, specifically when $p\geq\frac{3d-4}{d}$. This section leverages key tools such as the Banach-Alaoglu theorem (Subsection 4.3) and the Minty trick (Subsection 4.4). In Section 5, we address the decomposition of the pressure term, ensuring that each decomposed component corresponds to a distinct equation, as formalized in Theorem 5.1. The proof of Theorem 3.2 for small values of $p$, where $p\in\mathopen{}\mathclose{{\left(\frac{2d}{d+2},\frac{3d-4}{d}}}\right)$, is carried out in Section 6. This section introduces the regularized problem (6.3) and its corresponding definition (Definition 6.1). It also reconstructs the pressure term (Subsection 6.3) and employs compactness arguments (Subsection 6.4) alongside the monotonicity method (Subsection 6.5). Finally, the article concludes with Section 7, where we establish the uniqueness of the weak solution, as stated in Theorem 3.3.
@@ 2.2. Auxiliary results
$\mathrm{E}_{1}\hookrightarrow\hookrightarrow\mathrm{E}\hookrightarrow\mathrm{E}_{0}\quad\text{ and }\quad\|x\|_{\mathrm{E}}\leq C\|x\|_{\mathrm{E}_{0}}^{1-\theta}\|x\|_{\mathrm{E}_{1}}^{\theta},\text{ for any }\ x\in\mathrm{E}_{1}.$ (2.5)
@@
    $\mathrm{W}^{s_{0},p_{0}}(0,T;\mathrm{E}_{0})\cap\mathrm{W}^{s_{1},p_{1}}(0,T;\mathrm{E}_{1})\hookrightarrow\hookrightarrow\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad s<s_{\theta}\text{ and }s-\frac{d}{p}<s_{\theta}-\frac{d}{p_{\theta}};$
@@
    $\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad 0\leq s<1-\theta\text{ and }s-\frac{1}{p}<1-\theta-\frac{1}{p_{\theta}};$
@@ 3. Main results
$\bm{f}\in\mathbf{L}^{1}(Q_{T}),\qquad\bm{v}_{0}\in\mathbf{V},\qquad\frac{2d}{d+2}\leq p<\infty.$ (3.1)
@@
The main result of this work concerns the existence of a weak solution to problem (1.1). To prove this result, we strengthen the conditions in (3.1) on the forcing term, initial velocity, and range of the power-law exponent. In addition to the condition (3.5) imposed below, which strengthens the admissible range of the power-law exponent, we assume
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with a smooth boundary $\partial\Omega$ of class $\mathrm{C}^{2}$, and assume that hypotheses (3.3)-(3.4) are verified. If $2\leq d\leq 4$ and
$p>\frac{2d}{d+2},$ (3.5)
@@
Under the assumptions of Theorem 3.2, the solution to the problem (1.1) is unique in the class of weak solutions defined in Definition 3.1.
The proof of Theorem 3.2 is postponed to Sections 4 and 6, whereas the proof of Theorem 3.3 is carried out in Section 7.
Regarding the proof of Theorem 3.2, we proceed by distinguishing two separate cases:
$p\geq\frac{3d-4}{d}\quad\text{and}\quad\frac{2d}{d+2}<p<\frac{3d-4}{d}.$ (3.6)
This splitting is dictated by the following embedding, which refines the classical embedding (1.8) employed to establish the existence of weak solutions in [40]:
$\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})\hookrightarrow\mathrm{L}^{\rho}(0,T;\mathrm{L}^{\rho}(\Omega)^{d}),\quad\rho:=\frac{pd}{d-2},\quad d>2,$ (3.7)
which follows from the classical Sobolev embedding combined with parabolic interpolation arguments. The special cases $d=2$ and $d=p$ require less work, because the embeddings $\mathrm{W}^{1,2}_{0}(\Omega)^{d}\hookrightarrow\mathrm{L}^{\rho}(\Omega)^{d}$, when $d=2$, and $\mathrm{W}^{1,p}_{0}(\Omega)^{d}\hookrightarrow\mathrm{L}^{\rho}(\Omega)^{d}$, for $d=p$, are both valid for any $\rho\in(1,\infty)$. Under the embedding (3.7), one obtains
$t\longmapsto\int_{\Omega}\bm{v}(t)\otimes\bm{v}(t):\nabla\bm{\varphi}(t)\,\mathrm{d}\bm{x}\ \in\ \mathrm{L}^{1}(0,T),$ (3.8)
for all $\bm{v},\bm{\varphi}\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})\cap\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})$, provided that
$\frac{2}{\rho}+\frac{1}{p}\leq 1\quad\Leftrightarrow\quad p\geq\frac{3d-4}{d}.$
When $p$ satisfies the above bound, the space of test functions naturally coincides with the solution space, which simplifies the analysis considerably. In this case, the existence of weak solutions follows by standard application of the monotone operator theory [35, 40].
4. Existence: large values of $p$
In this section, we provide the proof of Theorem 3.2 in the case of
$p\geq\frac{3d-4}{d}.$ (4.1)
@@ 4.1. Approximate solutions
$\bm{v}_{0,n}\xrightarrow[n\to\infty]{}\bm{v}_{0},\quad\text{ in }\ \mathrm{W}^{1,2}_{0}(\Omega)^{d}\cap\mathbf{V}_{p},$ (4.7)
for $p$ satisfying the condition given in (4.1). Observe that the system (4.4)-(4.5) can be written in the matrix form as follows,
@@ 4.2. Uniform estimates
    If $\bm{v}_{0}\in\mathbf{V}$ and $p>1$, then there exists an independent of $n$ positive constant $C_{1}$ such that the following estimate holds true:
@@
    If $\bm{v}_{0}\in\mathbf{V}_{p}$, $2\leq d\leq 3$ and $p>1$, or $d>3$ and $p\geq d$, are verified, then there exists an independent of $n$ positive constant $C_{2}$ such that the following estimate holds true :
@@
To establish (4.10) and (4.11), we multiply equation (4.4) by $c_{k}^{n}(t)$ and $\frac{\mathrm{d}c_{k}^{n}(t)}{\mathrm{d}t}$, respectively, and sum the resulting expressions over $k=1$ to $k=n$. The desired estimates then follow from suitable applications of Hölder’s, Korn’s, Sobolev’s, Young’s, and Gronwall’s inequalities. The structure of the proof parallels the approach used in [7, Lemma 3.2], though our momentum equation differs from the one considered there. We just draw the reader’s attention to statement (2), where the distinction between the cases
$2\leq d\leq 3\quad\text{with}\quad p>1\qquad\text{and}\qquad d>3\quad\text{with}\quad p\geq d$
arises from the need to control the boundedness of the convective integral term
@@
When $d=2$ or $d=3$, the Ladyzhenskaya inequalities (cf. [35, Ch. 1, Lemmas 1–2]) can be applied as follows:
@@
$\mathopen{}\mathclose{{\left|\int_{\Omega}\big(\bm{v}_{n}(t)\cdot\nabla\big)\bm{v}_{n}(t)\cdot\partial_{t}\bm{v}_{n}(t)\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}(t)\|_{2}\,\|\nabla\bm{v}_{n}(t)\|_{p}\,\|\partial_{t}\bm{v}_{n}(t)\|_{\frac{2d}{d-2}},$
@@
$\frac{1}{2}+\frac{1}{p}+\frac{1}{\frac{2d}{d-2}}\leq 1\quad\Leftrightarrow\quad p\geq d.$
Remark 4.2.
Throughout this work, we focus on the parameter range (3.5), which constitutes the central contribution of this study. The restriction on the exponent $p$ appearing in statement (2) of Lemma 4.1, specifically the case $p\geq d$ when $d>3$, is of a technical nature and is mainly related to the boundedness of the convective term (4.12). For the purposes of the existence analysis, however, it is sufficient to obtain uniform bounds on the Galerkin approximated time derivative $\partial_{t}\bm{v}_{n}$ in $\mathrm{W}^{-s,2}(\Omega)^{d}$ in order to invoke the Aubin-Dubinskiĭ compactness lemma. Therefore, the precise restriction $p\geq d$ when $d>3$ is not essential for establishing the existence of weak solutions, it mainly facilitates the technical estimates for the nonlinear term. See Remarks 4.3 and 4.4 below.
@@
On the other hand, as a consequence of (3.7) and (4.10), we have
$\int_{0}^{T}\|\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)\|_{\frac{\rho}{2}}^{\frac{\rho}{2}}\,\mathrm{dt}\leq C,\qquad\rho:=\frac{pd}{d-2},$ (4.13)
@@
$\frac{\rho}{2}>1\Leftrightarrow p>\frac{2(d-2)}{d}\quad\Leftarrow\quad p\geq\frac{3d-4}{d}\ \text{ and }\ \frac{3d-4}{d}>\frac{2(d-2)}{d}\Leftrightarrow d>0.$
From (4.11), we have
$\int_{0}^{T}\|\partial_{t}\bm{v}_{n}(t)\|_{1,2}^{2}\,\mathrm{d}t\leq C.$ (4.14)
Using the embedding $\mathrm{W}_{0}^{1,2}(\Omega)^{d}\hookrightarrow\mathrm{L}^{q}(\Omega)^{d}$, which holds for $q=\frac{2d}{d-2}$ when $d\neq 2$, or for any finite $q\geq 1$ when $d=2$, together with the embedding
$\mathrm{W}_{0}^{1,r^{\prime}}(\Omega)^{d}\hookrightarrow\mathrm{L}^{q^{\prime}}(\Omega)^{d}\ \Leftrightarrow\ \mathrm{L}^{q}(\Omega)^{d}\hookrightarrow\mathrm{W}^{-1,r}(\Omega)^{d},$
which is valid under the condition
@@
$\mathopen{}\mathclose{{\left(\frac{2d}{d-2}}}\right)^{\prime}=q^{\prime}\leq(r^{\prime})^{\star}\ \Leftrightarrow\ r^{\prime}\geq\frac{2d}{d+4}\quad\Leftarrow\quad r^{\prime}>1,$
when $r^{\prime}<d$, or for any $r^{\prime}\in[1,\infty)$ when $r^{\prime}=d$, or any $r^{\prime}\in[1,\infty]$ when $r^{\prime}>d$, we deduce from (4.14) that
$\int_{0}^{T}\|\partial_{t}\bm{v}_{n}(t)\|_{-1,r}^{2}\,\mathrm{d}t\leq C,\qquad 1\leq r\leq\frac{2d}{d-2},$ (4.15)
for $d\neq 2$, or for any finite $r\geq 1$ when $d=2$.
Remark 4.3.
@@
Estimate (4.10) provides uniform bounds on $\bm{v}_{n}$ in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}(\Omega)^{d})$ and in $\mathrm{L}^{p}(0,T;\mathrm{W}^{1,p}(\Omega)^{d})$, independently of $n$. The Galerkin system (4.4) allows us to write
$(\mathrm{I}-\kappa\Delta)\partial_{t}\bm{v}_{n}=P^{n}\Big(-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n})+\nu\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))+\bm{f}\Big),$ (4.16)
@@
where $P^{n}$ is the orthogonal projection defined in (4.6). As $\kappa>0$, the elliptic operator $(\mathrm{I}-\kappa\Delta)$ is invertible, and, because $P^{n}$ acts on a finite-dimensional space, $(\mathrm{I}-\kappa\Delta)^{-1}P^{n}$ is bounded. Therefore it is sufficient to estimate the right-hand side of (4.16) in $\mathrm{W}^{-s,2}(\Omega)^{d}$ to control $\partial_{t}\bm{v}_{n}$.
@@
The convective term can be estimated, for any $\bm{\phi}\in\mathrm{W}_{0}^{s,2}(\Omega)^{d}$, as

### v2

@@ Abstract.
Here, for $d\in\{2,3,4\}$, $\bm{v}$ represents the velocity field, $\pi$ denotes the pressure, and $\bm{f}$ is the external forcing term. The constants $\kappa$ and $\nu$ correspond to the relaxation time and kinematic viscosity, respectively. The parameter $p\in(1,\infty)$ characterizes the fluid’s flow behavior, and $\mathbf{D}(\bm{v})$ denotes the symmetric part of the velocity gradient $\nabla\bm{v}$. For power-law exponents satisfying $p>1$ when $2\leq d\leq 3$, and $p>\frac{2d}{d+2}$ for $d=4$, we establish the existence of weak solutions to the generalised Navier-Stokes-Voigt system. Moreover, we prove uniqueness of the weak solution for the same ranges of $p$. The results are optimal in the sense that $p>1$ is minimal for $2\leq d\leq 3$. Moreover, for $p>\frac{2d}{d+2}$ with $d>3$, the framework uses a Gelfand triple, allowing the Aubin–Dubinskiĭ lemma to yield strong convergence of approximate solutions. This convergence is essential for the existence proof and holds precisely for $p>\frac{2d}{d+2}$ when $d=4$.
@@ 1.2. Literature review
$1<p<\frac{2d}{d+2},\quad\mbox{for}\ \ d>3,$
@@ 1.3. Objectives and novelties of the paper
  • As mentioned above, there are few works in the literature addressing power-law fluids [2, 3, 7, 21, 51]. In [51], the authors studied power-law fluids and established existence results for the exponent $p\in\big(\frac{2(d+1)}{d+2},\infty\big)$ using the $\mathrm{L}^{\infty}$-truncation, while [21] extended these results to $p\in\big(\frac{2d}{d+2},\infty\big)$ via Lipschitz-truncation. In our analysis, the presence of the Voigt term $\kappa\partial_{t}\Delta\bm{v}$ allows us to establish existence and uniqueness of the generalised Navier–Stokes–Voigt equations for the power-law exponents satisfying $p>1$ when $2\leq d\leq 3$, and $p>\frac{2d}{d+2}$ for $d=4$, using compactness and monotonicity arguments.
  • Another novelty of this article is that we do not employ the Helmholtz–Hodge projection, which would restrict our analysis to the case $p\in[2,\infty)$. Although the Helmholtz–Hodge projection is well-defined for $p\in(1,2)$, the analysis in this range is highly delicate and creates significant difficulties in controlling the convective term in the momentum equation (1.1)1. By avoiding the Helmholtz–Hodge projection, we are able to treat the full ranges $p>1$ for $2\leq d\leq 3$, and $p>\frac{2d}{d+2}$ for $d=4$, directly, without encountering these technical obstacles.
@@ 1.4. Organization
The structure of this article is organized as follows: Section 2 lays the groundwork by introducing the necessary function spaces and auxiliary results essential for the analysis. In Section 3, we present the definition of the weak solution (Definition 3.1) and state our main results, including Theorems 3.2, 3.3 and 3.4. Section 4 focuses on the proof of Theorem 3.2 for the cases of $p>1$ when $2\leq d\leq 3$, and $p\geq\frac{d}{2}$ when $d>3$. This section leverages key tools such as the Banach-Alaoglu theorem (Subsection 4.3) and the Minty trick (Subsection 4.4). In Section 5, we address the decomposition of the pressure term, ensuring that each decomposed component corresponds to a distinct equation, as formalized in Theorem 5.1. The proof of Theorem 3.3 for $\frac{2d}{d+2}<p<2$ when $d=4$, is carried out in Section 6. This section introduces the regularized problem (6.3) and its corresponding definition (Definition 6.1). It also reconstructs the pressure term (Subsection 6.3) and employs compactness arguments (Subsection 6.4) alongside the monotonicity method (Subsection 6.5). Finally, the article concludes with Section 7, where we establish the uniqueness of the weak solution, as stated in Theorem 3.4.
@@ 2.2. Auxiliary results
$\mathrm{E}_{1}\hookrightarrow\hookrightarrow\mathrm{E}\hookrightarrow\mathrm{E}_{0}\quad\wedge\quad\|x\|_{\mathrm{E}}\leq C\|x\|_{\mathrm{E}_{0}}^{1-\theta}\|x\|_{\mathrm{E}_{1}}^{\theta},\ x\ \in\mathrm{E}_{1}.$ (2.5)
@@
    $\mathrm{W}^{s_{0},p_{0}}(0,T;\mathrm{E}_{0})\cap\mathrm{W}^{s_{1},p_{1}}(0,T;\mathrm{E}_{1})\hookrightarrow\hookrightarrow\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad s<s_{\theta}\ \wedge\ s-\frac{d}{p}<s_{\theta}-\frac{d}{p_{\theta}};$
@@
    $\mathrm{W}^{s,p}(0,T;\mathrm{E}),\quad\mbox{if}\quad 0\leq s<1-\theta\ \wedge\ s-\frac{1}{p}<1-\theta-\frac{1}{p_{\theta}};$
@@ 3. Main results
$\bm{f}\in\mathbf{L}^{1}(Q_{T}),\qquad\bm{v}_{0}\in\mathbf{V},\qquad p>1.$ (3.1)
@@
The main result of this work concerns the existence of a weak solution to problem (1.1). For clarity of presentation, the existence result is divided into two theorems. The first corresponds to the case where the theory of monotone operators is sufficient, while the second addresses the case in which additional analytical tools are required.
For the first existence result, we strength conditions in (3.1) on the forcing term and initial velocity as follows
@@
$\bm{f}$ $\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-1,2}(\Omega)^{d})\cap\mathrm{L}^{p^{\prime}}(0,T;\mathrm{L}^{p^{\prime}}(\Omega)^{d}).$ (3.4)
Theorem 3.2.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with a Lipschitz-continuous boundary $\partial\Omega$, and assume that hypotheses (3.3)-(3.4) are verified. Suppose that either
$p>1\qquad\mbox{and}\qquad 2\leq d\leq 3,$ (3.5)
or
$p\geq\frac{d}{2}\qquad\mbox{and}\qquad d>3.$ (3.6)
For the second existence result, condition (3.4) can be relaxed to
@@
However, stronger assumptions on the regularity of the domain and on the spatial dimension will be required.
@@
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with a smooth boundary $\partial\Omega$ of class $\mathrm{C}^{2}$, and assume that hypotheses (3.3) and (3.7) are verified. If
$\frac{2d}{d+2}<p<\frac{d}{2}\qquad\mbox{and}\qquad d>3,$ (3.8)
@@
Observe that (3.8)1 is satisfied whenever $d>2$ and therefore, in particular, for $d>3$.
The proof of Theorem 3.3 will require the restriction $d\leq 4$ (see Section 6.4 and (6.65) below). Combined with the main assumption (3.8), this leaves only the case $d=4$. Consequently, the admissible range for the exponent $p$ reduces in fact to
$\frac{2d}{d+2}=\frac{4}{3}<p<2\qquad\mbox{and}\qquad d=4.$ (3.9)
Although physical applications are typically limited to $d\leq 3$, the four-dimensional case is analytically challenging. Many existence and regularity results for nonlinear PDEs critically depend on the interplay between the spatial dimension and the growth of the nonlinearity. Dimension $d=4$ represents a borderline dimension where the scaling of Sobolev embeddings and duality estimates is delicate. Establishing existence in this dimension shows the robustness of the analytical framework and provides insight into the mechanisms that control solutions in higher-dimensional settings. In this sense, Theorem 3.3 extends the applicability of the theory beyond physically relevant dimensions and highlights the structural stability of the approach under more demanding analytical conditions.
The next result addresses the uniqueness of weak solutions, showing that under the assumptions of Theorem 3.2, or Theorem 3.3, the problem (1.1) admits at most one solution in the sense of Definition 3.1.
@@
Under the assumptions of Theorems 3.2 and 3.3, the solution to the problem (1.1) is unique in the class of weak solutions defined in Definition 3.1.
The proofs of Theorems 3.2 and 3.3 are postponed to Sections 4 and 6, whereas the proof of Theorem 3.4 is carried out in Section 7.
The proofs of Theorems 3.2 and 3.3 are distinguished according to the cases (3.5)-(3.6) and (3.8). For Theorem 3.2, the result follows directly by applying the standard theory of monotone operators. In contrast, the proof of Theorem 3.3 requires, at the appropriate points, the introduction of technical tools and auxiliary results, which will be presented in the subsequent sections.
4. Proof of Theorem 3.2
The proof of Theorem 3.2 relies on the following embedding, which can be seen as a refinement of the classical embedding (1.8) used in the existence theory of weak solutions in [40]:
$\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})\cap\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})\hookrightarrow\mathrm{L}^{\rho}(0,T;\mathrm{L}^{\rho}(\Omega)^{d}),$ (4.1)
where
$\rho:=\max\mathopen{}\mathclose{{\left\{\frac{2d}{d-2},\frac{dp}{d-p}}}\right\}.$ (4.2)
This embedding follows from the classical Sobolev inequality combined with standard parabolic interpolation arguments. The value of $\rho$ defined in (4.2) is mainly relevant in the case $d>\max\{2,p\}$. Indeed, when $d=2$ or $d=p$, the exponent $\rho$ can take arbitrarily large values in $1,\infty)$, while if $d<p$ it is even possible to set $\rho=\infty$. This reflects the fact that in these borderline or low-dimensional cases the embedding (4.1) can be made as strong as needed, which greatly simplifies the control of the nonlinear terms.
On the other hand, the information provided solely by $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})$ is sufficient to show that
$t\longmapsto\int_{\Omega}\bm{v}(t)\otimes\bm{v}(t):\nabla\bm{\varphi}(t)\,\mathrm{d}x\ \in\ \mathrm{L}^{1}(0,T),$ (4.3)
for all $\bm{v},\bm{\varphi}\in\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})$. The key point is that (4.3) does not impose any restriction on the exponent $p$, and thus this argument applies for all $p>1$. Moreover, (4.3) plays a central role in controlling the convective term, which allows the direct application of the theory of monotone operators to establish existence, in the spirit of [35, 40].
In the following, we present the proof of Theorem 3.2, which is divided into several subsections. The proof is included for the sake of completeness and because some of its arguments will be used later in the proof of Theorem 3.3.
@@ 4.1. Approximate solutions
$\bm{v}_{0,n}\xrightarrow[n\to\infty]{}\bm{v}_{0},\quad\text{ in }\ \mathrm{W}^{1,2}_{0}(\Omega)^{d}\cap\mathbf{V}_{p}.$ (4.9)
Observe that the system (4.6)-(4.7) can be written in the matrix form as follows,
@@ 4.2. Uniform estimates
    If $\bm{v}_{0}\in\mathbf{V}$ and $p>1$, then there exists a positive constant $C_{1}$, independent of $n$, such that the following estimate holds true:
@@
    Let $\bm{v}_{0}\in\mathbf{V}_{p}$, and assume that either hypothesis (3.5) or (3.6) is satisfied. Then, there exists a positive constant $C_{2}$, independent of $n$, such that the following estimate holds:
@@
To establish (4.12) and (4.13), we multiply equation (4.6) by $c_{k}^{n}(t)$ and $\frac{\mathrm{d}c_{k}^{n}(t)}{\mathrm{d}t}$, respectively, and sum the resulting expressions over $k=1$ to $k=n$. The desired estimates then follow from suitable applications of Hölder’s, Korn’s, Sobolev’s, Young’s, and Gronwall’s inequalities. The structure of the proof parallels the approach used in [7, Lemma 3.2], though our momentum equation differs from the one considered there. We just draw the reader’s attention to statement (2), where the distinction between the cases (3.5) or (3.6), arises from the need to control the boundedness of the convective integral term
@@
When $d=2$ or $d=3$, the Ladyzhenskaya inequalities (cf. [35, Chapter 1, Lemmas 1–2]) can be applied as follows:
@@
$\mathopen{}\mathclose{{\left|\int_{\Omega}\big(\bm{v}_{n}(t)\cdot\nabla\big)\bm{v}_{n}(t)\cdot\partial_{t}\bm{v}_{n}(t)\,\mathrm{d}x}}\right|\leq\|\bm{v}_{n}(t)\|_{\frac{2d}{d-2}}\,\|\nabla\bm{v}_{n}(t)\|_{p}\,\|\partial_{t}\bm{v}_{n}(t)\|_{\frac{2d}{d-2}},$
@@
$\frac{2}{\frac{2d}{d-2}}+\frac{1}{p}\leq 1\quad\Leftrightarrow\quad p\geq\frac{d}{2}.$
@@
On the other hand, as a consequence of (4.1)-(4.2) and (4.12), we have
$\int_{0}^{T}\|\bm{v}_{n}(t)\otimes\bm{v}_{n}(t)\|_{\frac{\rho}{2}}^{\frac{\rho}{2}}\,\mathrm{dt}\leq C,\qquad\rho=\max\mathopen{}\mathclose{{\left\{\frac{2d}{d-2},\frac{dp}{d-p}}}\right\},$ (4.15)
@@
$\rho=\mathopen{}\mathclose{{\left\{\begin{array}[]{ll}\displaystyle\frac{2d}{d-2},&1<p\leq 2,\\[10.0pt] \displaystyle\frac{dp}{d-p},&p\geq 2,\end{array}}}\right$.
and thus
@@
$\frac{\rho}{2}>1\Leftrightarrow\mathopen{}\mathclose{{\left\{\begin{array}[]{ll}\displaystyle\frac{2d}{d-2}>2,&1<p\leq 2,\\[10.0pt] \displaystyle p>\frac{2d}{d+2},&p\geq 2.\end{array}}}\right$.
Therefore, the uniform estimate (4.15), together with the fact that $\frac{\rho}{2}>1$, which will be used later (see (4.29) below), does not impose any restriction on $p$.
We now show that
$\mbox{$\partial_{t}\bm{v}_{n}$ \ is uniformly bounded in \ $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d})$},$ (4.16)
@@
for $s$ satisfying (4.4). In fact, estimates (4.12) and (4.13) provide uniform bounds on $\bm{v}_{n}$ in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,2}_{0}(\Omega)^{d})$ and in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{1,p}_{0}(\Omega)^{d})$, independently of $n$. The Galerkin system (4.6) allows us to write
$(\mathrm{I}-\kappa\Delta)\partial_{t}\bm{v}_{n}=P^{n}\Big(-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n})+\nu\,\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))+\bm{f}\Big),$
@@
where $P^{n}$ is the orthogonal projection defined in (4.8). As $\kappa>0$, the elliptic operator $(\mathrm{I}-\kappa\Delta)$ is invertible, and, because $P^{n}$ acts on a finite-dimensional space, $(\mathrm{I}-\kappa\Delta)^{-1}P^{n}$ is bounded. Therefore it is sufficient to estimate
$\Big(-\mathbf{div}\,(\bm{v}_{n}\otimes\bm{v}_{n})+\nu\,\mathbf{div}\,(|\mathbf{D}(\bm{v}_{n})|^{p-2}\mathbf{D}(\bm{v}_{n}))+\bm{f}\Big)$ (4.17)
@@
in $\mathrm{L}^{\infty}(0,T;\mathrm{W}^{-s,2}(\Omega)^{d})$ to control $\partial_{t}\bm{v}_{n}$. The convective term can be estimated, for any $\bm{\phi}\in\mathrm{W}_{0}^{s,2}(\Omega)^{d}$, as